OpenStudy (anonymous):

Sharmi ran 3 miles 24 in minutes. How long would it take her to run 7 miles?

hartnn (hartnn):

3 miles 24 minutes, so, 1 mile = 24/3 =.... ?

OpenStudy (anonymous):

56 minutes

OpenStudy (anonymous):

if 24/3 is dividing the answer would be 8

hartnn (hartnn):

thats for 1 minute. for 7 minutes, it'll be 7*8 =...? minutes

OpenStudy (anonymous):

56 minutes

hartnn (hartnn):

that would be correct.

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome ^_^